regarding question #201706, how did you get the value of x?
Answer by Alan3354(69443)

find angle A if it supplements angle B which is equal to 70-x
-------------
A + B = 180
A + (70-x) = 180
A - x = 110
A = x + 110
---------
x can be any value (if you allow negative angles).
If, eg, x = 10, then
A = 120, B = 60
if x = 15,
A = 125, B = 55